OpenStudy (anonymous):

What is the coordinate of the center of the circle given by the equation (x-11)^2+(y+22)^2=1? A. (1,2) B. (11,-22) C. (-11,22) D. (121,484)

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

The general form of a circle is given by:\[\Large\rm (x-h)^2+(y-k)^2=r^2\]With radius \(\Large\rm r\) and center \(\Large\rm (h,k)\)

zepdrix (zepdrix):

See how the general form has subtraction next to each x and y? We need that subtraction so we can accurately determine the coordinate of the center.

OpenStudy (anonymous):

ok

zepdrix (zepdrix):

\[\Large\rm (x-11)^2+(y+22)^2=1\]So if we want subtracting with our y.... we can maybe write it like, minus a negative, that's the same as a positive,\[\Large\rm (x-11)^2+(y-[-22])^2=1\]

zepdrix (zepdrix):

I hope it's not too confusing doing it that way >.< Maybe an easier way too remember would be... you always want the opposite sign of whatever you see in the brackets. Since it's x-11, our x coordinate of the center will be 11.

zepdrix (zepdrix):

Since it's y+22, our y coordinate will be -22.

zepdrix (zepdrix):

That make... some sense...? >.< heh

OpenStudy (anonymous):

yes it makes complete since your awesome! thank you!

OpenStudy (anonymous):

sense*

zepdrix (zepdrix):

lol np :3